Find median for the following data:

X 5 10 15 20 25

f 25 20 15 10 5


1
Expert's answer
2021-01-13T19:46:05-0500

To find the median we first calculate a table which is given below



Now "N=\sum f_i=75"

Therefore median "M=(\frac{N+1}{2})"th term

"=(\frac{75+1}{2})"th term

"=38" th term

"=10"